9.09 points Following is partial information for Delamunte Industries for the month of August Balance, August 1 Direct materials used Direct labor Manufacturing overhead applied Process S 41,000 75,000 120,000 90,000 Jobs finished during August are summarized here: of Jobs Job # 234 237 231 246 Completed S 58,000 5,000 74,500 67,500 Required At the end of August, only one job, Job 248, was stl in process. The direct labor cost incurred on Job 248 as of August 31 was $24,000 Calculate the predetermined overheed rate that was used to apply manufacturing overhead to jobs during August. (Round your answer to 2 decimal places) per Direct Labor dola

Required: At the end of August, only one job, Job 248, was still in process. The direct labor cost incurred on Job 248 as of August 31 was $24,000 1. Calculate the predetermined overhead rate that was used to apply manufacturing overhead to jobs during August. (Round your answer to 2 decimal places.) per Direct Labor dollar 2. Determine the balance in Work in Process at the end of August 3. Determine how much of the direct materials cost and applied ovorhead would be recorded tor Job 248 as of August 31. (Round your Predetermined Overhead Rate to 2 decimal places.) Direct Materials Cost Applied Overhead
Answer #1

1. Predetermined Overhead Rate = 90000/120000 = $0.75 per direct labour.

2. Balance in work in process = Opening Work in Process + Manufacturing Cost - Cost of Goods Manufactured

= 41000 + (75000+ 120000 + 90000) - (58000 + 65000 + 74500 + 67500) = 41000 + 285000 - 265000 = $61,000

3. It is given that only job 248 was in process as on August 31. From above we found out that the cost of job 248 was 61000.

Overhead applied rate is 0.75 per direct labour dollar and direct labour cost incurred in job 248 is 24000. So the applied manufacturing overhead amounts to $18,000 (24000*0.75)

So, the Direct Material Cost will amount to $19,000 (61000 - 24000 - 18000).